OpenStudy (anonymous):

Let logb 2 = 2.218 and logb 3 = 3.413. Find logb 24.

OpenStudy (anonymous):

log24=log(8*3)=log8+log3=3log2+log3=3(2.218)+3.413=10.067

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome